OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -2, and -1 + 2i here are my choices f(x) = x^4 - 7x^2 - 26x - 40 f(x) = x^4 - 3x^3 - 8x^2 - 13x - 40 f(x) = x^4 + 6.5x^2 - 26x - 40 f(x) = x^4 - 3x^3 + 8x^2 + 13x + 40

OpenStudy (anonymous):

this again ready?

OpenStudy (anonymous):

i dont know how to do this at all lolol

OpenStudy (anonymous):

ok then we better go slow

OpenStudy (anonymous):

do you know if \(r\) is a zero, then one factor must be \((x-r)\)? so in this case you have \(4\) is a zero making one factor \((x-4)\)

OpenStudy (anonymous):

clear or not? that is really what we need to start this

OpenStudy (anonymous):

oh okay so (x-4) and (x+2) ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

we start with \[(x-4)(x+2)\] now we have more work to do

OpenStudy (anonymous):

you also need the quadratic that has the zeros of \(-1+2i\) and it "conjugate" \(-1-2i\) next job is to find that

OpenStudy (anonymous):

any ideas? "no" is a fine answer

OpenStudy (anonymous):

no i dont know

OpenStudy (anonymous):

ok there is a hard way, an easy way, and a really really easy way lets do the easy way first, then the real easy way

OpenStudy (anonymous):

work backwards starting with \[x=-1+2i\] add \(1\) to get \[x+1=2i\]

OpenStudy (anonymous):

then square both sides (carefully) to get \[(x+1)^2=(2i)^2\] or \[x^2+2x+1=-4\]

OpenStudy (anonymous):

add 4 to both sides to finish with \[x^2+2x+5\]

OpenStudy (anonymous):

final job is to multiply \[(x-4)(x+2)(x^2+2x+5)\]i would cheat so as not to screw up the algebra

OpenStudy (anonymous):

and would that be the answer? or is there more

OpenStudy (anonymous):

OOO got it. thank you

OpenStudy (anonymous):

yeah that is it

OpenStudy (anonymous):

yw